OpenStudy (anonymous):

solve for the varible in this equation... x/2-2/3=1/4

OpenStudy (johnweldon1993):

You would add 2/3 to both sides...to isolate where 'x' is on the left remember you cannot add fractions without a common denominator then divide both sides by 2 \[\frac{ x }{ 2 }-\frac{ 2 }{ 3 }=\frac{ 1 }{ 4 }\] after adding 2/3 to both sides \[\frac{ x }{ 2 }=\frac{ 1 }{ 4 }+\frac{ 2 }{ 3 }\] so the least common denominator of (1/4) and (2/3) is?

OpenStudy (anonymous):

um, 12? maybe?

OpenStudy (johnweldon1993):

right 12 so we can do that out now \[\frac{ 3 }{ 12 } + \frac{ 8 }{ 12 } = \frac{ 11 }{ 12 }\] so now we have \[\frac{ x }{ 2 } = \frac{ 11 }{ 12 }\] so now we multiply both fractions by 2 to completely isolate 'x'

OpenStudy (anonymous):

okay. so how would you do that?

OpenStudy (anonymous):

WELL IT JUST DEPENDS

OpenStudy (anonymous):

on what?

OpenStudy (johnweldon1993):

\[2*\frac{ x }{ 2 } = x\] \[2 * \frac{ 11 }{ 12 } = \frac{ 11 * 2 }{ 12 }\] \[\frac{ 22 }{ 12 }\] so you have \[x = \frac{ 22 }{ 12 }\]

OpenStudy (anonymous):

Ohhhh. So whats the answer? I et what your saying but, im very confused on how its typed out.

OpenStudy (johnweldon1993):

that is the answer 22/12 or if you would prefer a decimal 1.83333333333 Try and plug it into the equation for 'x'

OpenStudy (johnweldon1993):

It's an ugly answer I know, but sometimes that how math problems go...

OpenStudy (anonymous):

okay, thanks.

OpenStudy (johnweldon1993):

no problem...I hope you understand it though?

OpenStudy (texaschic101):

once you get x/2 = 11/12....this is a proportion and you cross multiply (12)(x) = (11)(2) 12x = 22 (now divide by 12 to get x by itself) 12x/12 = 22/12 x = 22/12 reduces to 11/6